Two small nonconducting spheres have a total charge of Q=Q1+Q2= 90.0 μC , Q1<Q2. When placed 31.0 cm apart, the force each exerts on the other is 13.0 N and is repulsive.

Part A: What is the charge Q1?

Part B: What is the charge Q2?

Part C: What would Q1 be if the force were attractive?

Part D: What would Q2 be if the force were attractive?
Answer #1

Given,

Q1 + Q2 = 90

Q2 = 90 - Q1

Force, F = KQ1 Q2/r^2

13 = (9 x 10^9) Q1 x (90 - Q1) * 10^-12 / 0.31^2

- Q1^2 + 90 Q1 - 138.81 = 0

Q1 = 1.57 uC

Q2 = 88.43 uC

C)

Q1 = 90 + Q2

13 = (9 x 10^9) ( 90 + Q2) Q2 *10^-12 / 0.31^2

Q2^2 + 90 Q2 - 1 38.81 = 0

Q1 = 1.52 uC

Q2 = - 91.52 uC

Comment in case any doubt please rate my answer....